# Evil twins are not that evil: Qualitative insights into machine-generated prompts

### v1

@@ 4.2 Replacing autoprompt tokens
Working from now on with the pruned autoprompts, we replace the token in each position in turn with one of the 10k most frequent tokens from the Pile. We quantify the impact of the ablations in terms of BLEU score with respect to the original continuation. The ablation results are summarized in Fig. 3, where replacements are binned based on the impact they have on the continuation (examples are presented in the tables of Appendix A).
Figure 3: Average proportions of replacement effect types by position on pruned autoprompts, aligned from right (whiskers show standard deviations). Null-effect replacements leave the continuation unchanged. Moderate replacements have BLEU of at least 0.2. Strong replacements have BLEU below 0.2.
First, we confirm that non-pruned tokens in all positions play a significant role in generating the continuation, as shown by the fact that most replacements have a strong impact on BLEU. However, for all positions except the last, we also see that a non-negligible proportion of replacements do not affect the continuation at all, and in a significant proportion of cases the continuation is only mildly affected (as the examples in Table 8 of Appendix A show, even a BLEU score around 0.2 typically corresponds to a continuation that is quite similar to the original).
We confirm moreover the special role of the last token, that can almost never be replaced without a catastrophic result on the continuation. The importance of the ending of the autoprompt is further shown by the fact that, as we approach the last position, it is increasingly more difficult to find replacements that do not strongly affect the continuation.
Furthermore, by manually inspecting the cases that lead to only a moderate change in the continuation, we observe that sometimes they show a degree of “compositionality”, in the sense that the continuation stays the same except for one or a few tokens that are replaced with new tokens that reflect the meaning of the replacement, and/or drift away from the meaning of the replaced token. Some examples are presented in Table 3.
To make this intuition more quantitative, we ran the following experiment. First, to facilitate automated similarity analysis, we extracted all cases where the replacement leads to the change of a single (typographic) word in the continuation (about 3% of the total cases). For these cases, we used FastText (Bojanowski et al., 2017) to measure the semantic similarity of both the original autoprompt token and its replacement to the original word in the continuation and to the changed one. We found that the original token is more similar to the new continuation word (vs. the original one) in only 37% of the cases, whereas the replacement token is more similar to the new continuation in 55% of the cases. We thus conclude that, indeed, there is a tendency for at least this type of replacement to work compositionally, with a small change in the autoprompt leading to a semantically consistent change in the continuation. This, in turn, suggests that autoprompts do not function as unanalyzable holistic wholes, but their “meaning” to the model derives, at least partially, from assembling the meaning of its parts, as with natural language sequences. As the examples show, though, this assembling looks nothing like the one performed by natural language syntax.
autoprompt: continuation:
cake implies Norman meaning LOVE/radical journalism indicated by the use of the word "love/radical" in the title.
sDad/meal ———— Protection Many mans ruggedally understands the need to protect his family/food
@@
Table 3: Example autoprompt token replacements leading to a small, interpretable change in the continuation (replaced/replacement tokens in the autoprompt and changed material in the continuation are highlighted in bold).
@@ 4.3 Shuffling autoprompt tokens
The picture we get from the previous studies is one where autoprompts are composed of three types of tokens. A number of tokens are fillers that, being ignored by the LM, can simply be pruned. The final token is extremely important and hard to change, because, in autoregressive prediction, it determines the exact nature of the first token of the continuation, and consequently strongly affects the rest of the continuation. The other non-prunable tokens also have an impact on the continuation, but they seem to rather work as single “keywords” that affect the semantic content of what follows, without forming a tight syntactic bound with each other and what follows.
Previous work has uncovered a somewhat mixed picture in terms of the robustness of autoprompts to token order shuffling (Ishibashi et al., 2023; Rakotonirina et al., 2023; Melamed et al., 2024). Based on what we just observe, we conjecture that the last token will be “rigid”, as moving it around would strongly affect the continuation, whereas the preceding tokens might be more robust to order ablations. To test the conjecture, we randomly shuffled the tokens (10 repetitions per autoprompt) and measured the resulting BLEU with respect to the original continuation. We either shuffled all tokens or left the last one fixed.
The average BLEU when shuffling all tokens is at 0.02 (s.d. 0.03) and at 0.05 (s.d. 0.07) when leaving the last token in its slot. This difference is highly significant (paired t-test, $p<0.001$).The difference stays comparably significant if, in the first condition, we leave a random non-last token fixed, so that the same number of tokens is shuffled in the two cases. However, the low BLEU values suggest that, contrary to our conjecture, the autoprompt tokens before the last are not a bag of keywords, since their order matters as well. One possibility is that, while autoprompts as a whole do not constitute syntactically well-formed sequences, they are composed of tight sub-sequences that should not be separated. For example, given that modern tokenizers split text at the sub-word level, token-level shuffling will arbitrarily break words.
Some support for the view that the catastrophic effect of shuffling pre-last tokens is due to short-distance dependencies comes by looking at the cases in which a bigram in an autoprompt (excluding the last position) is also attested in the Pile corpus, either in the original or in the inverted order. In 61.5% of these cases, the Pile frequency of the original bigram is larger than that of the inverted one. This suggests that there is at least some tendency towards a natural local ordering among autoprompt tokens.
@@ 4.4 Making human prompts more autoprompt-like
Applying the same greedy-pruning method to the original prompts, we find that more than 99% can also be pruned, with 21.9 tokens removed on average. Considering the average token length of the original prompts is 38.57, this means that, strikingly, on average 57% of the tokens can be removed without affecting the continuation. Since the prompts are long, one could think that what is removed is primarily material towards the beginning of the sequence, but actually we find that 95% of the prompts also have pruned tokens among the last 10 items.
Examples of the latter are in Table 4. Prunable material often consists of modifiers whose removal does not affect the basic syntactic structure of the fragments (“strategic bomber”, “section of the pipeline”, “replication fork”…), but this is not always the case, and in many examples pruning turns well-formed sentences into seemingly unstructured token lists or telegraphic texts at best ( “most section, since it”, “fork mobile day but”). Still, like in the case of the autoprompts, the coherence of the transition between the prompt and the continuation is generally preserved (“…bomber Tu, / which was designed…”, “…since it / was the only one…”).
prompt: continuation:
… Soviet prototype strategic bomber based on the Tu 4, which was designed to replace the Tu-4.
… most complex single section of the pipeline, since it was the only one that was not under contract with Fluor.
formers (16–18 year olds) were recruited for the performance.
replication fork and the mobile Holliday junction, but the structure of the DNA duplex was not known.
… the 74 gun Theseus, provided an escort and fired a salvo of shells at the enemy’s batteries.
Table 4: Randomly selected examples of original prompts with prunable tokens in bold. Only the last 10 tokens of each original prompts are shown. In the first example, the last token of the autoprompt is a comma, which is not pruned. In the third example, the brackets are not pruned, either.
Table 5 shows the original-prompt tokens that are most typically kept vs. pruned. As for the autoprompts (cf. Table 2 above), the highly prunable tokens consist entirely of common function words and punctuation marks. However, the typically kept tokens tend to also consist of (somewhat rarer) function words and punctuation marks. The presence of quotation marks and brackets in this list should not surprise us, because removing these elements from the prompt will strongly affect the continuation (e.g., without the opening bracket in the prompt, the model might fail to close a parenthetical, producing a completely different continuation). However, what is the crucial distinction between the typically kept vs. pruned function words is not clear to us, and it deserves further investigation in the future.
kept: pruned:
" the
which of
was a
is in
) ’s
when on
film and
with The
Table 5: Top-10 kept (left) or pruned (right) original prompt tokens ranked by local mutual information.
Figure 4 presents pruning proportion by position for the last 10 tokens in the original prompts, confirming that, in this case as well, the last token is by far the most important one in determining the continuation. Interestingly, the contrast is even more dramatic than for autoprompts (cf. Figure 1 above).
Figure 4: Counts of original prompt tokens that were pruned (dark orange) and kept (yellow) in the last 10 positions, where 0 is the last position.
@@
We replicate the token-replacement experiment on the pruned original prompts, obtaining the results summarized in Figure 5, where we used the same BLEU ranges as in Figure 3 above. Again, tokens become more replaceable as we move away from the end of the prompt, confirming the crucial role played by the very last token.
Figure 5: Average proportions of replacement effect types by position on pruned original prompts, aligned from right, limited to the last 10 tokens (whiskers show standard deviations).
Table 6 show examples in which the original prompt, despite pruning and replacement among the last 10 tokens, still triggers the same continuation. We see how the same principles that might explain the success of autoprompts are at work here, suggesting how autoprompts might take shape during their induction process. For example, both “…citing the “popular adventure book”, attempted the first” and “… adventure regression first” trigger the continuation “ascent of the north face of Mount Everest.” The last token is preserved, and determines the fact that the continuation will start with a noun. The term adventure probably contributes to determine that the continuation is something adventurous, but, as the materials surrounding the token have been deleted, it acts more like a keyword than a proper syntactic element. Finally, the irrelevant inserted token regression is ostensibly ignored.
original: … one of the best examples of American surrealism and
modified: … one of003 and
continuation: one of the best films of the 1990s.
original: … I Ever Wanted, and “Already Gone
modified: … Ifold, and “Alreadyone
continuation: ” was the first single released from the album.
original: … citing the “popular adventure book”, attempted the first
modified: adventure regression first
continuation: ascent of the north face of Mount Everest.
Table 6: Examples where pruning and replacing a token in an original prompt does not affect the continuation. The original row shows the last 10 tokens of an original prompt; the modified row shows the equivalent prompt suffix after pruning and replacement.
@@
Shuffling all tokens of the original prompts after pruning leads to an average BLEU of 0.02 (s.d. 0.03), comparable to what observed for autoprompts. Leaving the last token in place leads to an average BLEU of 0.03 (s.d. 0.05). This small difference is again highly significant (paired t-test, $p<0.001$), confirming the importance of the last token for the subsequent prediction (the difference stays equally significant if we compare shuffling all but the last token to shuffling while keeping one random non-last token fixed).
@@ 5 Discussion
We show that seemingly opaque, machine-induced prompts possess, to some extent, interpretable properties, such as a strong reliance on the last token, the presence of filler tokens that are ignored by the model, and the compositional-like behavior of some keyword tokens. We further observe that some of these properties are also present in natural prompts.
These findings might shed some light on how LMs process language in general. They seem to rely on a simplified model of it, where not all tokens have specific syntactic and semantic functions in an abstract syntactic tree. We note that the phenomenon of relying on over-simplified representations of the data is not specific to LMs. Convolutional Neural Network classifiers of visual data have also been shown to latch onto superficial correlations in the data, leading to poor out-of-distribution generalization (Jo and Bengio, 2017; Ilyas et al., 2019; Yin et al., 2019; Geirhos et al., 2020).
Identifying and characterizing the features that deep learning models respond to are crucial steps in understanding their inner workings and making them more robust. In future work, besides addressing the issues discussed in the Limitations section below, we aim to extend our analysis beyond discrete tokens, focusing either on circuits through mechanistic interpretability methods or on representations using a more top-down approach, such as representation engineering (Zou et al., 2023a).
Limitations
  • Due to the time it takes to induce autoprompts with our computational resources, we could only experiment with 3 models, the largest of which has 6.9B parameters. We make our code available in hope that researchers with bigger resources will run similar experiments on a larger scale.
  • For analogous reasons, we only experimented with one variant of the autoprompt inducing algorithm, and we fixed the number of tokens in the induced prompt to 10. Given that all algorithms we are aware of adopt similar gradient-following methods, and based on qualitative inspections of autoprompt examples in other papers, we expect our conclusions to hold for autoprompts independently of how they are induced, but this should be verified empirically.
  • Our autoprompts most closely resemble adversarial attack where an obfuscated sequence is used to retrieve one specific piece of information from the LM. However, autoprompts might be also induced for other purposes, such as to improve factual knowledge retrieval when combined with a query sequence (Shin et al., 2020). It remains to be explored if different classes of autoprompts possess significantly different properties.
  • We have now a basic understanding of how an autoprompt determines its continuation, but we still need a better characterization of which tokens are more likely to be pruned, and of the means by which randomizing non-last tokens affects the continuation so strongly.
@@
If we do not achieve a genuine understanding of how LMs process and generate text, we cannot fully control their behaviour and mitigate unintended or intentional harm. Opaque autoprompts are an indication that there are important aspects of LM prompting and generation that are still out of our control. Our investigation into the nature of this phenomenon contributes to a better understanding of how LMs work and, thus, ultimately, to make them safer and more predictable.

### v3

@@ 4.2 Replacing autoprompt tokens
(a) Autoprompts
(b) Original prompts (last 10 tokens)
Figure 4: Average proportions of replacement effect types by position on pruned prompts, aligned from right (whiskers show standard deviations). Null-effect replacements leave the continuation unchanged. Moderate replacements have BLEU $\geq 0.2$. Strong replacements have BLEU $<0.2$.
(a) Autoprompts
(b) Original Prompts
Figure 5: Semantic similarity between substitutes. When a token admits fewer substitutes, they generally lie in a smaller semantic space than when the substitute count is large. The rightmost data points correspond to the most flexible tokens and tend to be semantically inconsistent. The leftmost data points are tokens admitting 0 or few substitutes, which are semantically closer.
Methodology
Working from now on with the pruned autoprompts, we replace the token in each position in turn with one of the 10k most frequent tokens from the Pile. We quantify the impact of the ablations in terms of BLEU score with respect to the original continuation. The ablation results are summarized in Fig. 4(a), where replacements are binned based on the impact they have on the continuation (examples are presented in the tables of Appendix B).
Impact of the replacements
First, we confirm that non-pruned tokens in all positions play a significant role in generating the continuation, as shown by the fact that most replacements have a strong impact on BLEU. However, for all positions except the last, we also see that a non-negligible proportion of replacements do not affect the continuation at all, and in a significant proportion of cases the continuation is only mildly affected (as the examples in Table 11 of Appendix B show, even a BLEU score around 0.2 typically corresponds to a continuation that is quite similar to the original). We confirm moreover the special role of the last token, that can almost never be replaced without catastrophic results. More generally, as we approach the last position, it is increasingly more difficult to find replacements that do not strongly affect the continuation.
Autoprompts Generated Continuation
WHM modelling tag Mus before either forming militant Annex Ireland …or joining the Provisional IRA
@@
forming: ’ meeting’, ’ leaving’, ’ remaining’, ’ developing’, ’ selling’, ’ breaking’, ’ producing’, …
Ireland: ’ Irish’
either/ militant: 0 substitutes.
<|endoftext|> <|endoftext|> Star Defense *]{} 1950 blamed Tehran instead rhe …United States for the 1953 coup.
[…]
Tehran: ’ Iran’
instead: ’ than’, ’=’, ’ then’, ’ back’, ’ to’, ’&’, ’ again’, ’ though’, …
blamed/ rhe: 0 substitutes.
Original Prompt Generated Continuation
for television Services Our is create programming that offers …a unique experience for our viewers.
[…]
create: ’ about’,’ emotional’,’Create’,’about’,’ differentiation’,’ unusual’,’ demands’,’ Create’,…
that: ’that’
offers: ’ provide’,’ provides’,’ offer’
programming: 0 substitutes.
Doctor can transcend reach ep iph any and greater …understanding of the human condition.
[…]
iph: ’ph’,’rep’
any: ’ understanding’,’ening’,’aining’,’ insight’
and: ’ for’,’ by’,’ or’,’A’,’ &’,’ through’,’or’,’an’,’ upon’,’ via’,’OR’,’ toward’,’ towards’,’ AND’,’ gain’, …
greater: ’ improved’,’ enhanced’,’ deeper’,’ wider’
Table 4: Examples of replacement for autoprompts and original prompts (10 last tokens). Color represents the number of substitutes: $>50$ substitutes and $<50$ substitutes. When there are less than 50 substitutes, a random subset is displayed. Difficult-to-render characters are replaced by ’?’. More examples are provided in Table B (appendix).
Fillers and key tokens
We further looked at the equivalent sets, consisting, for a given token, of the substitutes that keep the continuation unchanged. Table 3 shows $76.4\%$ of the tokens can be substituted by at least one equivalent, and each token has $302.8$ substitutes on average. But the size of the equivalent set is highly variable, with half of the tokens having $15.0$ or fewer equivalents. This disparity is intuitively informative about the role of each token. In particular, we identify the presence of fillers that can be replaced by a large number of substitutes, and key tokens, that admit more restrained equivalent sets and must thus carry more specific information. This is evidenced by the fact that language-like tokens tend to admit less equivalent substitutes than non-linguistic tokens. Table 4.2 (top section) shows examples of autoprompts with tokens color-coded based on their equivalent set size. When the equivalent set is small, we provide a random sample of it as an illustration.
Semantic consistency of the equivalent sets
We measured the semantic consistency within equivalent sets in order to determine whether substitution in autoprompt is governed by semantic similarity (akin to synonymy in natural language) or by other factors. Specifically, we used FastText (Bojanowski et al., 2017) to compute the average semantic similarity: (a) between a token and each of its equivalents; and (b) among the equivalents themselves (including the original token). [^9] For technical reasons, due to the difficulty of obtaining reliable representations for non-linguistic tokens, we focused only on language-like tokens. The first measure indicates how semantically relevant the substitutes are in relation to the original token, while the second measure reflects the size of the substitute semantic space. A value close to $1$ suggests that the substitutes cluster within a small semantic region, meaning they are semantically consistent with one another. We plot semantic similarity against set size in Fig 5(a). As expected, small equivalent sets –corresponding to key tokens– tend to be semantically coherent on average. In this sense, they approximate near synonyms in natural language. This is confirmed by the examples in Table 4.2, e.g., ‘Ireland’ is replaced by ‘Irish’ and ‘Tehran’ by ‘Iran’. Nonetheless, the semantic relation is often approximate, e.g., ‘before’ being replaced by ‘then’, ‘change’ by ‘subsequently’ or ‘forming’ by ‘meeting’, ‘leaving’, ‘remaining’, etc. In contrast, large sets –associated to fillers– often include tokens that are semantically unrelated to each other. We show in Table 5 a subset of the most frequently found tokens in large sets. It mostly consists of subwords (sometimes interpretable morphemes), digits and fragments of named entities that do not carry strong semantic information.
105 <EOT> () 113 ini Mar ran 25 be Green 101 56 ¸ mat ato ator Rock 35 115 Moore 114 SS lim eter ester abe uns Head Char ? ? ? ac aa aur ? ong ost den FF 54 a ab ud sl rim st ? 022 1 37 91 INT op ore ef rel cha ast ives
Table 5: List of the $60$ tokens most frequently found in large equivalent sets (observed in about $10\%$ of the sets). Difficult-to-render characters are replaced by ’?’. Tokens are marked as follows: digit, artifact and punctuation, named entity, subword and morphological unit. See more in Table 14.
Can humans predict which autoprompt tokens are more important?
The four authors of the paper manually annotated the autoprompt dataset with a binary label marking, for each token, if it is intuitively important or not for the generation of the continuation. The annotation was made a posteriori, using the full autoprompt and the generated continuation (see appendix C for a description of the annotation process). Results show that autoprompts possess interpretable properties, as the labeling is correlated with the number of replacements a token might have. Indeed, we measured that the median size of the equivalent set for tokens deemed important is $2$, against $15$ for those deemed replaceable, with the effect observed both for language-like and non-linguistic tokens (cf. Table 16).
Traces of compositionality
In cases where the replacement causes only a moderate change in the continuation, we see signs of “compositionality,” in the sense that the continuation only displays a few new tokens broadly reflecting the meaning of the replacement. Some examples are presented in Table 6. We make the intuition more quantitative as follows. First, to facilitate automated similarity analysis, we extracted all cases where the replacement leads to the change of a single (typographic) word in the continuation (about 3% of the total). For these cases, we used FastText to measure the semantic similarity of both the original autoprompt token and its replacement to the original word in the continuation and to the changed one. We found that the original token is more similar to the new continuation word (vs. the original one) in only 48% of the cases, whereas the replacement token is more similar to the new continuation in 59% of the cases. We thus conclude that, indeed, there is a tendency for at least this type of replacement to work compositionally, with a small change in the autoprompt leading to a semantically consistent change in the continuation. This, in turn, suggests that autoprompts do not function as unanalyzable holistic wholes. Their “meaning” to the model derives, at least partially, from assembling the meaning of its parts, as with natural language sequences. However, this looks nothing like the one performed by natural language syntax.
Autoprompts Generated Continuation
Eg<EOT> Brown mushrooms/face chooses suffix "brown …" to describe the color of the mushroom/face.
crossesFootnote Several panels accidentally have feather/bands 517 chant collectively …, as if they were a single bird/band.
fecture Phoenix/Toronto Latinoamous]], effectively …making it the largest city in Arizona/Canada.
@@
Table 6: Example autoprompt token replacements leading to a small, interpretable change in the continuation (legend: replaced/replacement in autoprompt; and original/new in continuation).
@@ 4.3 Shuffling autoprompt tokens
Methodology
Previous work has uncovered a somewhat mixed picture in terms of the robustness of autoprompts to token order shuffling (Ishibashi et al., 2023; Rakotonirina et al., 2023; Melamed et al., 2024). Based on our ad-interim observations, we conjecture that the last token will be “rigid”, as moving it around would strongly affect the continuation, whereas the preceding tokens might be more robust to order ablations. To test the conjecture, we randomly shuffled tokens (10 repetitions per autoprompt) and measured the resulting BLEU with respect to the original continuation. We either shuffled all tokens or left the last one fixed.
More than a bag-of-words
The average BLEU when shuffling all tokens is at 0.03 (s.d. 0.04) and at 0.06 (s.d. 0.11) when leaving the last token in its slot. This difference is highly significant (paired t-test, $p<0.001$). [^10] The difference stays comparably significant if, in the first condition, we leave a random non-last token fixed, so that the same number of tokens is shuffled in the two cases. However, the low BLEU values suggest that, contrary to our conjecture, the autoprompt tokens before the last are not a bag of keywords, since their order matters as well. One possibility is that, while autoprompts as a whole do not constitute syntactically well-formed sequences, they are composed of tight sub-sequences that should not be separated. For example, given that modern tokenizers split text at the sub-word level, token-level shuffling will arbitrarily break words. Some support for the view that the catastrophic effect of shuffling pre-last tokens is due to short-distance dependencies comes by looking at the cases in which a bigram in an autoprompt (excluding the last position) is also attested in the Pile corpus, either in the original or in the inverted order. In 60.5% of these cases, the Pile frequency of the original bigram is larger than that of the inverted one, suggests some degree of natural local ordering among autoprompt tokens.
@@ 4.4 Making human prompts more autoprompt-like
Applying the same greedy-pruning method to the original prompts, we find that more than 99.5% can also be pruned, with 23.8 tokens removed on average (s.d. 13.2). Considering the average token length of the original prompts is 39.0, this means that, strikingly, on average 61% of the tokens can be removed without affecting the continuation. Since the prompts are long, one could think that what is removed is primarily material towards the beginning of the sequence, but actually we find that 95% of the prompts also have pruned tokens among the last 10 items. Examples of the latter are in Table 1 (bottom). Prunable material often consists of modifiers whose removal does not affect the basic syntactic structure of the fragments (“causing the body’s immune cells”, “Aviation Regiment”), but this is not always the case, and in many examples pruning turns well-formed sentences into seemingly unstructured token lists or telegraphic text at best. Still, like in the case of the autoprompts, the coherence of the transition between the prompt and the continuation is generally preserved (“…a third / of which…”, “…as her / parents are…”). Table 2 (right) shows the original-prompt tokens that are most typically kept vs. pruned. As for the autoprompts, highly prunable tokens consist entirely of common function words and punctuation marks. However, typically kept tokens might also be (somewhat rarer) function words and punctuation marks. Figure 2(b) presents pruning proportion by position for the last 10 tokens in the original prompts, confirming that, in this case as well, the last token is by far the most important one in determining the continuation. Interestingly, the contrast is even more dramatic than for autoprompts (Figure 2(a)).
@@
We replicate the token-replacement experiment on the pruned original prompts, obtaining the results summarized in Figure 4(b). Again, tokens become more replaceable as we move away from the end of the prompt, confirming the crucial role played by the very last token. Moreover, we confirm the presence of both key tokens, having few semantically related substitutes, and fillers with numerous semantically inconsistent substitutes (Figure 5(b)). Table 4.2 (bottom) shows examples in which the original prompt, despite pruning and replacement among the last 10 tokens, still triggers the same continuation. We see how the same principles that might explain the success of autoprompts are at work here, suggesting how autoprompts might take shape during the induction process.
@@
Shuffling all tokens of the original prompts after pruning leads to an average BLEU of 0.02 (s.d. 0.02), comparably to autoprompts. Leaving the last token in place leads to an average BLEU of 0.03 (s.d. 0.05). This small difference is highly significant (paired t-test, $p<0.001$), confirming the importance of the last token (the difference stays equally significant if we compare shuffling all but the last token to shuffling while keeping one random non-last token fixed).
@@ 5 Discussion
Our findings about autoprompts, confirmed by autoprompt-inspired ablations of natural prompts, suggest that LMs might rely on a simplified model of language, where not all tokens have specific syntactic and semantic functions in an abstract syntactic tree. We note that the phenomenon of relying on over-simplified representations of the data is not specific to LMs. Convolutional Neural Network classifiers of visual data also latch onto superficial correlations in the data, leading to poor ood generalization (Jo & Bengio, 2017; Ilyas et al., 2019; Yin et al., 2019; Geirhos et al., 2020).
While we hope our results are of general interest, we recognize a number of limitations. First, due to the time it takes to induce autoprompts with our computational resources, we could only experiment with 6 models, the largest of which has 12B parameters. We make our code available in hope that researchers with bigger resources will run similar experiments on a larger scale. For analogous reasons, we only experimented with one variant of the autoprompt inducing algorithm. Given that all algorithms we are aware of adopt similar gradient-following methods, and based on qualitative inspection of autoprompt examples in other papers, we expect our conclusions to hold for autoprompts independently of how they are induced, but this should be verified empirically.
@@
If we do not achieve a genuine understanding of how LMs process and generate text, we cannot fully control their behavior and mitigate unintended or intentional harm. Opaque autoprompts are an indication that there are important aspects of LM prompting and generation that are still out of our control. Our investigation into the nature of this phenomenon contributes to a better understanding of how LMs work and, thus, ultimately, to make them safer and more predictable.